### v2

# ECLARE: Efficient cross-planar learning for anisotropic resolution enhancement

Abstract
In clinical imaging, magnetic resonance (MR) image volumes are often acquired as stacks of 2D slices with decreased scan times, improved signal-to-noise ratio, and image contrasts unique to 2D MR pulse sequences. While this is sufficient for clinical evaluation, automated algorithms designed for 3D analysis perform poorly on multi-slice 2D MR volumes, especially those with thick slices and gaps between slices. Super-resolution (SR) methods aim to address this problem, but previous methods do not address all of the following: slice profile shape estimation, slice gap, domain shift, and non-integer or arbitrary upsampling factors. In this paper, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a self-SR method that addresses each of these factors. ECLARE uses a slice profile estimated from the multi-slice 2D MR volume, trains a network to learn the mapping from low-resolution to high-resolution in-plane patches from the same volume, and performs SR with anti-aliasing. We compared ECLARE to cubic B-spline interpolation, SMORE, and other contemporary SR methods. We used realistic and representative simulations so that quantitative performance against ground truth can be computed, and ECLARE outperformed all other methods in both signal recovery and downstream tasks. Importantly, as ECLARE does not use external training data it cannot suffer from domain shift between training and testing. Our code is open-source and available at https://www.github.com/sremedios/eclare.
1 Introduction
Magnetic resonance (MR) imaging (MRI) is an essential clinical tool in diagnosis, monitoring, management, and research across many health conditions. Multi-slice 2D MRI uses slice selection [1] to acquire multiple evenly spaced 2D slices, which are stacked to create a 3D volume. In contrast, 3D MRI uses two phase encode directions to acquire 3D volumes directly [2]. Although 3D MRI has many advantages, including greater utility in downstream processing, multi-slice 2D MRI still accounts for the majority of clinical scans because it: 1) is faster to acquire; 2) offers clinically relevant tissue contrasts difficult to acquire in 3D [3]; 3) has relaxed hardware requirements; 4) is very familiar to clinicians; and 5) can be less sensitive to motion artifacts. For these reasons, multi-slice 2D MRI are common in pragmatic clinical trials [4] and in pre-clinical imaging studies [5].
Multi-slice 2D MR neuroimages can be strikingly anisotropic. Although in-plane resolutions for 2D MRI are frequently less than 1 mm in both phase-encode and readout directions, slice thicknesses may be 3–8 mm and many acquisitions additionally have slice gaps of 0.5–2 mm. Automated processing pipelines, which are often designed and trained for isotropic image volumes, perform sub-optimally on such anisotropic volumes (even when interpolated to isotropic voxel spacing). Super-resolution (SR) aims to solve this problem by estimating high-resolution (HR) images from low-resolution (LR) acquisitions. Most contemporary work on SR for MRI focuses on deep neural networks. Convolutional neural networks (CNNs) form the backbone of many approaches [6, 7], including generative adversarial networks (GANs) [8, 9, 10], normalizing flows [11, 12, 13, 14], and denoising diffusion probabilistic models (DDPMs) [15, 16, 17, 18, 19, 20]. The transformer architecture [21, 22, 23, 24, 25] is also common, sometimes used alongside CNNs [26, 27, 28, 29].
Figure 1: A flowchart of our proposed method. The anisotropic input volume is fed independently into each of the three steps. First, in (A) (Sec. 3.2), we estimate the slice excitation profile with ESPRESO [30]. Second, in (B) (Sec. 3.4), we extract HR in-plane 2D patches and use the PSF estimated from (A) to create paired training data. This training data is used to train the network $f_{\theta}$ with supervised learning. Third, in (C) (Sec. 3.5), we extract LR through-plane 2D slices and super-resolve them with the trained network $f_{\theta}$ from (B).The super-resolved slices are stacked and averaged, yielding the super-resolved output volume.
Although many SR approaches demonstrate remarkable performance, none at present satisfy all desirable properties for SR of multi-slice 2D MRI. Most methods, for example, do not consider the slice profile and true through-plane resolution of 2D MRI. Indeed, many approaches conflate slice separation with through-plane resolution, but they are different concepts and must be handled differently. For example, methods evaluated on simulations that use LR data that is only downsampled from 3D volumes (i.e., with interpolation) and without proper simulation of slice profile and slice separation would be more properly labeled as “inpainting” or “slice imputation” rather than super-resolution. As well, many methods in the literature have been trained for super-resolution of single integer factors such as 2 and 4 [6, 31, 11, 15, 21, 26]. Not only do these methods require retraining for each integer factor that is encountered, they also use techniques like power-of-two upsampling or pixel shuffle [32] that do not allow for non-integer factors. Finally, many methods are trained on specific MR tissue contrasts and often (in neuroimaging) on skull-stripped data, necessitating retraining and possibly re-evaluation for alternative tissue contrasts and for with-skull data. At present, there is no algorithm that handles all of these concerns for an arbitrary multi-slice 2D MR image volume.
To address these concerns, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a deep learning algorithm to super-resolve multi-slice 2D MR volumes along the through-plane direction. ECLARE builds on a succession of self-SR approaches [33, 34, 35, 36], in which all training data is extracted from the HR in-plane slices within a LR image volume. However, ECLARE adds several crucial features to properly model slice acquisition, account for non-integer super-resolution factors, and incorporate efficient network design elements. In particular, we use ESPRESO [30], a GAN-based algorithm previously developed by our team, to estimate the relative slice selection profile. We use this profile to model slice thickness and gap and to simulate training data from HR in-plane patches extracted from the anisotropic input volume. These data are used to train an efficient SR neural network architecture based on “wide activation SR” (WDSR) [37]. The model is then applied to both canonical through-planes, yielding two super-resolved volumes that are then averaged to create the final output. Underlying all interpolation operations in our approach is a novel field-of-view (FOV)-aware interpolation algorithm that respects the image coordinate system and target sample spacing. The main contributions of this paper are the following:
  • We propose an FOV-aware interpolation to resample an image to a new grid while preserving a desired sample spacing and maintaining the location of the FOV.
  • We model the relationship between isotropic and anisotropic volumes with explicit slice thickness and slice gap, then leverage ESPRESO to learn this model.
  • We propose an efficient alternative to previous self-SR approaches in MR: a single-model architecture without pre-trained weights to perform both anti-aliasing and SR.
  • We propose an architectural change to WDSR to yield the correct number of slices at the correct FOV at both integer and non-integer anisotropy factors.
  • We evaluate our approach and compare against several methods in both signal recovery metrics and downstream task performance in two disparate datasets.
Our proposed method super-resolves an anisotropic volume without external training data and can serve as an alternative to spatial interpolation with improved quality.
2 Related Work
In this work, we consider SR for 2D multi-slice MRI. While there is a body of literature discussing SR approaches focusing on training a model with external data, we do not discuss it here. Such approaches create paired LR-HR data through simulation from large datasets, perform supervised learning with or without consideration of arbitrary scale factors, and either must be applied to a similar domain as the training set or undergo domain adaptation for practical use. Alternatively, in this paper we consider methods that aim to be applicable “out-of-the-box.” Accordingly, our discussion of related work is limited to self-supervised or foundational models for SR of 2D anisotropic MR volumes. Thus, at test-time, the existence of relevant external training data is not assumed.
Figure 2: A 1D demonstration of the differences between interpolation implementations. The input has 6 pixels and its FOV center is located at 2.5, indicated by the red dashed line. Its sample spacing is 1, indicated by the purple arrow. Its extents occur at -0.5 and 5.5, indicated by the blue dashed line. In this demonstration, the desired sample spacing for upsampling is 0.7, and the desired sample spacing for downsampling is 1.429. The implementation scipy.zoom [38] maintains the FOV center and boundary sample locations, but changes the sample spacing to 0.625 and 1.667 for upsampling and downsampling, respectively. The implementation torch.nn.functional.interpolate [39] anchors the FOV extent to start at 0.5, shifting the center from the desired FOV center. Our proposed method (Sec. 3.3) maintains both the sample spacing and the FOV center.
Self-SR in MRI was first explored [33] by fitting an anchored neighborhood regression [40] model to learn the mapping from LR to HR patches, followed by Fourier burst accumulation (FBA) [41] to aggregate estimated high frequency Fourier coefficients into the super-resolved volume. The related work SMORE [34] reinterpreted self-SR using CNNs, which greatly increased the output accuracy and visual appearance. SMORE trains two enhanced deep residual networks (EDSRs) [42]: one for anti-aliasing and one for SR. These networks are applied in cascade to orthogonal through-plane slices, and all intermediate results are fused with FBA. Furthermore, although SMORE can be a self-SR method, [34] found that pre-training on similar data (i.e., human head MR images) and fine-tuning on the anisotropic volume at test-time allowed fewer training iterations and achieved better results.
Implicit neural representations (INR) [43, 44, 45] learn the mapping from coordinate to intensity. Since coordinates can be sampled at any resolution, this has inspired several works related to arbitrary-scale SR [46, 45, 47]. Recently, multi-contrast INR [48] was proposed to super-resolve subjects with multiple available contrasts from the same session, ideally acquired at different orientations (e.g., sagittal $T_{1}$-w and coronal T2-w). However, that method also supports the use of a single contrast, training in a similar manner to self-SR methods.
SynthSR [49] is a joint synthesis and SR model. Training data for SynthSR includes real MPRAGE head volumes and their corresponding parcellations. Domain randomization uses a Gaussian mixture model to produce synthetic MR-like images of various contrasts alongside augmentations such as simulated anisotropy, slice separations, bias fields, and additive noise. Through domain randomization, SynthSR maps an MR image of arbitrary contrast to a 1 mm isotropic T1-w MPRAGE scan.
3 Methods
Our proposed ECLARE improves resolution along the through-plane axis of an anisotropic 2D-acquired multi-slice MR image volume. To explain ECLARE, we first model the relationship between HR and LR volumes (Sec. 3.1). We then describe how to approximate the slice selection profile (Sec. 3.2) and FOV-aware interpolation (Sec. 3.3). We approximate the slice selection profile with ESPRESO [30] and use it to train a CNN with in-plane HR patches and their simulated LR counterparts (Sec. 3.4). Finally, we show how ECLARE uses this 2D CNN to super-resolve the 3D volume (Sec. 3.5) and describe implementation details (Sec. 3.6).
Figure 3: A diagram of our proposed architecture, based on WDSR [37]. The low resolution image $y$ is first processed by a convolution to produce a feature map with 256 channels. It then passes through several WDSR blocks. A final convolution is done to produce the correct number of channels for pixel shuffle. Finally, the result is downsampled and added to an upsampled $y$ to yield the output. In our implementation, we use 16 WDSR blocks.
3.1 Modeling the imaging process
A 2D-acquired MR image volume is a stack of independent 2D slices. In-plane, data is acquired in k-space and transformed to the image domain through the inverse Fourier transform. For through-plane, the relationship between slices is entirely characterized in the image domain. Common slice profiles take the shape of truncated Gaussians or use the Shinnar-Le Roux algorithm [50, 51] to approximate a rect. All in-plane measurements are subject to the same slice thickness and location.
We denote the slice selection profile as $h$. It is a continuous function of the through-plane position. We note that although slice selection is implemented as a bandpass filter, relative to a particular through-plane axis position the slice profile is a low-pass filter. This corresponds to the observation that ideal slice profiles are rect-like or Gaussian-like.
An ideal multi-slice 2D MR volume $x$ would be acquired in an isotropic fashion with resolution $a\times a\times a$ mm. For practical reasons (time and physical constraints), however, a real multi-slice 2D MR volume $y$ is acquired with thicker slices—i.e., $a\times a\times b$ mm where $b>a$. A further problem is that $y$ may have slice gaps where the slice separation in the through-plane direction is greater than $b$, the slice thickness. We model 2D multi-slice acquisitions as a discrete convolution between the unobserved HR image volume and the slice selection profile:
$y=(x\ast h)\downarrow_{r},$ (1)
where $\downarrow_{r}$ is the subsampling ratio of $r$. The full-width-at-half-max (FWHM) of $h$ determines the slice thickness and $r$ is a ratio between the HR slice separation $s^{\ast}$ and the LR slice separation $s$, defined as $r=\frac{s}{s^{\ast}}$. If $h$ has a smaller FWHM than the separation $s$, then Eq. 1 models slice gap. Similarly, a FWHM of $h$ equal to or greater than $s$ respectively models adjacent slices and slice overlap.
3.2 Approximating the imaging process
The imaging process described by Eq. 1 comprises blurring with $h$ and subsampling by $r$. The subsampling factor $r$ is known since the HR and LR slice separations are each known, so the remaining task is to determine the slice selection profile $h$. In general, $h$ is a product of a real-world implementation that is subject to imperfections in the instrumentation, and hence cannot be derived precisely. However, since our intent is to use HR in-plane patches as the target distribution, estimating $h$ directly is not necessary. Instead, we leverage ESPRESO [30] to estimate a relative slice selection profile $\tilde{h}$, shown in Fig. 1 (A). The desired relative slice selection profile $\tilde{h}$ blurs data from the in-plane slices such that the distribution matches the LR through-plane data. To estimate $\tilde{h}$, ESPRESO optimizes a neural network to produce a fixed-size discrete profile through adversarial training. It trains on a single LR anisotropic volume by assuming that degraded in-plane and through-plane patches are indistinguishable in distribution for an appropriate patch size. Briefly, training consists of extracting in-plane image patches, degrading by $\tilde{h}$, comparing to through-plane patches, and backpropagating the loss through a discriminator. The optimal $\tilde{h}$ yields degraded in-plane image patches that are indistinguishable in distribution from true through-plane image patches. We use this estimated $\tilde{h}$ in lieu of $h$ for Eq. 1 hereafter.
3.3 FOV-aware interpolation
While interpolation is routine in digital signal processing, the exact implementation has several freedoms. ECLARE requires both downsampling (for training data simulation) and upsampling (for inference) and we would like both to be stable with respect to the FOV. We consider three conditions that are desirable to meet with respect to the FOV of the signal:
  1. 1.
    the center of the FOV should be at the same location;
  2. 2.
    the samples should be spaced as specified;
  3. 3.
    the locations of boundary samples should be maintained.
Interpolation algorithms cannot meet all of these conditions in general since only a discrete number of samples is permitted. However, since it is possible to satisfy two of these conditions, a compromise on the third is necessary. We quickly review popular modern implementations. The method scipy.ndimage.zoom [38] meets conditions 1) and 3). It adjusts the sample spacing to preserve the FOV center while locking the locations of boundary samples. The method torch.nn.functional.interpolate [39] meets only condition 2). It preserves the sample spacing, but neither the FOV center nor the boundary sample locations. Instead, it anchors the extent on the first sample. We place more importance on the center of the FOV and the sample spacing and therefore desire conditions 1) and 2). We summarize the interpolation method proposed in [52] that maintains the FOV center and sample spacing while sacrificing the precise location of the boundary samples. A comparison of the three implementations is shown in Fig. 2.
Figure 4: A $5\|1.5$ sagittal slice from a representative subject from the OASIS-3 dataset. The first row shows the image domain with PSNR and SSIM for the entire image volume overlaid. The second row shows the simplified SLANT labels, with an overlay showing the mean cDSC across all labels for the entire image volume.
As in [52], we introduce a fixed coordinate system for the signal before and after interpolation. Let the position of the first pixel be centered at zero, let $N$ be the number of pixels before interpolation, and let the spacing before interpolation be one. The width of each pixel is equal to the spacing; before interpolation it is therefore also one. The first pixel $p_{0}$ is centered at zero and the last pixel $p_{N-1}$ is centered at $N-1$. Since each pixel has a width of one, $p_{0}$ extends from $-0.5$ to $0.5$ and $p_{N-1}$ extends from $N-1.5$ to $N-0.5$. Therefore the entire image extends from $-0.5$ to $N-0.5$.
[…]
$N^{\prime}=\left\lfloor\frac{N}{d}\right\rceil.$ (2)
Accordingly, we notate the first pixel after interpolation as $p^{\prime}_{0}$ and the last pixel as $p^{\prime}_{N^{\prime}-1}$.
In order to center around the same FOV, the center position of the entire image should be identical before and after interpolation. Since the center of the FOV is the halfway point between the first and last pixels, we can write
$c=\frac{1}{2}(p_{0}+p_{N-1})$ $=\frac{1}{2}(p^{\prime}_{0}+p^{\prime}_{N^{\prime}-1})$ (3)
Also, since we know the new spacing $d$ and the new number of pixels $N^{\prime}$, we can calculate the position of any pixel $p_{i}^{\prime}$, $i=0,1,2,\ldots,N^{\prime}-1$ in terms of the new starting pixel:
$p^{\prime}_{i}=p^{\prime}_{0}+di.$ (4)
To find the position of the first pixel $p^{\prime}_{0}$, we use the center $c$:
$c$ $=\frac{1}{2}(p^{\prime}_{0}+p^{\prime}_{N-1})$ (5)
$=\frac{1}{2}(p^{\prime}_{0}+[p^{\prime}_{0}+d(N^{\prime}-1)])$ (6)
$p^{\prime}_{0}$ $=-\frac{d}{2}(N^{\prime}-1)+c.$ (7)
Finally, we can calculate the new extents of the entire image after interpolation. The first pixel extends from $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{0}+\frac{d}{2}$ and the last pixel extends from $p^{\prime}_{N-1}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$. Thus, the image extents are $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$.
The result of this method yields an FOV-aware interpolation implementation that maintains the FOV center and guarantees the desired pixel spacing. This is true in both upsampling and downsampling scenarios as demonstrated in Fig. 2.
3.4 Training a deep learning model with paired data
We now develop a relationship between in-plane and through-plane data and train a SR network. Paired LR and HR data are necessary for supervised learning. We extract HR 2D patches from in-plane and produce blurred 2D patches using Eq. 1 with the learned $\tilde{h}$ from ESPRESO. These blurred 2D patches are then downsampled using the FOV-aware interpolation described in Sec. 3.3. Thus, each LR patch has the appropriate simulated slice thickness, separation, and FOV.
A deep convolutional neural network model $f_{\theta}$ learns the mapping between the paired patches, as shown in Fig. 1 (B). We use an architecture based on WDSR [37] after empirically evaluating other architectures such as EDSR [42] and RCAN [31]. The WDSR architecture, designed for efficiency, achieves the same quantitative metrics as other architectures but trains in less time. However, the WDSR architecture itself uses pixel shuffle [32], which requires integer upsampling factors, to perform the final upsampling to the desired image size. In general, though, the required upsampling ratio $r$ may be non-integer. To address this, we rewrite $r$ as
$r=\lceil r\rceil\frac{r}{\lceil r\rceil},$ (8)
[…]
3.5 Inference
The trained network is fully convolutional, allowing it to be used on full-sized image slices despite training on patches. Our inference strategy to super-resolve the volume is to independently super-resolve all slices from each canonical through-plane axis, stack the super-resolved slices into two volumes, and take an unweighted mean of the two volumes to produce the final output. This procedure is illustrated in Fig. 1 (C).
Figure 5: Quantitative results (PSNR on the left, SSIM on the right) on the OASIS-3 dataset across resolutions.
3.6 Implementation details
ECLARE is implemented in PyTorch and we train the model from scratch with each new multi-slice 2D MR image. We optimize the parameters of the network by minimizing mean squared error using a total of $1,000,000$ patches with a batch size of $128$ patches. The training patches are of size $\lceil 8r\rfloor\times 8$. We use a $21$-tap filter to represent $h$, which follows the standard implementation of ESPRESO. We used the Adam [53] optimizer, a learning rate of $10^{-3}$, the one-cycle learning rate scheduler [54], and PyTorch’s automatic mixed precision feature. Our implementation uses $16$ residual blocks of $256$ channels. With these hyperparameter choices, on an NVIDIA V100 the total running time, including self-training and inference, is just under 5 minutes. The memory usage is 1.16 GB on the GPU. Our code is open-source and is available at https://www.github.com/sremedios/eclare.
4 Experiments and Results
4.1 Data and evaluation metrics
Our experiments began with HR isotropic 3D-acquired image volumes of resolution 1×1×1 $1\times 1\times 1\text{~{}mm}^{3}$. These images were then degraded with Eq. 1, setting $h$ to the shape of the slice excitation profile created by the Shinnar-Le Roux (SLR) pulse design algorithm [50, 51], $s$ to the desired slice spacing (described below), and $x$ to a 3D-acquired isotropic volume. The implementation of the SLR algorithm we use is from the sigpy library [55].
[…]
We compare ECLARE to bicubic interpolation, SynthSR [49], single-contrast INR [48], and SMORE [34]. Since its output is always a T1-w MPRAGE, we do not evaluate SynthSR on the FLAIR SR task. Other mainstream SR methods are not appropriate for comparison here due to the heterogeneities in data acquisition strategies for multi-slice 2D MRI. Many LR volumes have non-integer ratios to achieve isotropic resolution or have slice gap. Furthermore, there is a need to evaluate on arbitrary image contrasts without retraining for each subtle change in contrast. ECLARE and the methods we compare to are agnostic to these differences.
Since we have ground truth (GT) images for simulated data, we can use reference-based image quality metrics to evaluate SR performance. We use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) with the implementations provided by the scikit-image Python library [58]. We additionally evaluate the utility of the super-resolved images by segmentation downstream tasks. We use the consistency Dice similarity coefficient (cDSC), which we define as the Dice similarity coefficient (DSC) between an automated algorithm’s performance on the HR image volume and the super-resolved image volume. This is to differentiate cDSC from DSC which commonly is used between an algorithm-generated label and a human-generated label.
Figure 6: Quantitative results on cDSC by anatomical region on the OASIS3 dataset across resolutions. In the interest of space, only four representative regions (namely, ventricles, cerebral white matter, caudate, and cerebellar white matter) are shown.
4.2 Healthy T1-w simulations
We calculated cDSC on the SLANT [59] and SynthSeg [60] estimates of the OASIS-3 dataset. Since SLANT produces many small labels, we simplify them down to $9$ labels as in our previous work [36]: ventricles, cerebellar gray matter (GM), cerebellar white matter (WM), cerebral GM, cerebral WM, caudate, putamen, thalamus, and brainstem. SynthSeg is a generalized automated segmentation technique designed to be robust to a variety of image appearances. It directly produces volume estimates of similar regions of interest, so we include the following comparable regions: ventricles, cerebral white matter, cerebral gray matter, cerebellar white matter, cerebellar grey matter, thalamus, putamen, caudate, and brainstem.
On the OASIS-3 dataset, we show qualitative results from a representative subject at $5\|1.5$ resolution in Fig. 4. The image domain is shown alongside the SLANT result labels across the methods. In the image domain, it is apparent that our proposed ECLARE method is visually closer to the ground truth (GT) in both appearance and contrast. Viewing Fig. 4 column-by-column, in the image domain bicubic interpolation does not address the extreme aliasing, but is relatively accurate with respect to the SLANT labels. SynthSR generates a realistic brain and head appearance, but mismatches the contrast and hallucinates some structures such as the cortical sulci that can be seen in the SLANT labels row. Single-contrast INR does not generate accurate results, synthesizing a different brain stem appearance from the other methods, as shown in the green brainstem label in the second row. The SMORE results highlight an “over-sharpening” effect with hyperintense regions near the skull and superior to the ventricles and the bottom part of the brainstem. ECLARE yields the most visually accurate image of the group in both the image as well as qualitative inspection of the segmentation. The SLANT segmentation results are still not ideal for any method, and some areas of cortical gray matter are missed by all methods.
We show the PSNR and SSIM for the five methods in Fig. 5. Across all resolutions, our proposed ECLARE outperforms all other methods in both metrics except for the resolution $3\|0$, where SMORE is best.
In Fig. 6, we show the cDSC for the nine regions of interest calculated by SLANT [59]. Overall we see similar trends as in Fig. 5, with ECLARE performing best at all resolutions except at $3\|0$, where SMORE performs best. An exception to this is the cerebellar white matter, where at nearly every resolution SMORE is most performant. An interesting observation is how robust SLANT is to interpolation artifacts; bicubic interpolation outperforms single-contrast INR and SynthSR consistently across all resolutions and structures.
Figure 7: Radar plots showing the $R^{2}$-value for a linear regression of all $N=50$ subjects between SynthSeg volume estimation on the super-resolution method vs. the SynthSeg volume estimation of the ground truth.
We also used SynthSeg [60] as a downstream task. For all subjects, we fit a linear regression model to the sets of estimated volumes against the GT volumes per region per resolution where the dependent variable was the GT volume and the independent variable was the SR volume. The fit of the line is quantified by the $R^{2}$ coefficient of determination, which we plot in Fig. 7. Some regions (ventricles, cerebral white matter, brainstem) are consistently well-segmented at all resolutions across all SR methods, but others show clear winners. ECLARE outperforms all methods consistently, visualized by the enclosure of other methods on the radar plots.
4.3 People with MS: T2-w FLAIR simulations
One prominent characteristic of MS is the presence of pathology as white matter hyperintense lesions. Data-driven SR algorithms conventionally must include relevant pathology in their training data to be expected to perform well on pathological data. ECLARE, SMORE, and single-contrast INR accomplish this intrinsically since they train on in-plane slices, many of which contain lesions. For the 3D-MR-MS dataset, we used the automated white matter lesion segmentation algorithm SELF [61] to produce labels for the GT and super-resolved volumes, comparing results with cDSC. We also included signal recovery metrics PSNR and SSIM as with the previous dataset.
Figure 8: A $4\|1.2$ mid-sagittal slice for a representative subject from the 3D-MR-MS dataset. Each column shows the MR image with the automated lesion segmentation overlaid in neon-green. The PSNR, SSIM, and cDSC for the entire image volume of this representative are overlaid.
We performed similar analyses for the 3D-MR-MS dataset, but with SynthSR omitted due to the image being T2-w FLAIR. Qualitative results are shown in Fig. 8, where the automated segmentation by SELF is overlaid in neon-green. This T2-w FLAIR clearly shows the “over-sharpening” artifact from SMORE in the hyperintense region inferior to the ventricle. Many such examples of hyperintense regions just inferior to a dark surface are visible in the SMORE result, whereas the ECLARE result does not exhibit this. Furthermore, ECLARE better recovers from aliasing artifacts induced by the gap, since this image is $4\|1.2$.
Figure 9: Quantitative results (PSNR in the top-left, cDSC in the top-right, SSIM in the bottom-left, and a zoomed y-axis of SSIM in the bottom-right) on the 3D-MR-MS dataset across resolutions. The bottom-right zoomed SSIM is to better illustrate comparisons between methods with more accurate SSIM scores.
In Fig. 9, we show the PSNR and SSIM for signal recovery and cDSC for lesion segmentation consistency. Like in the OASIS-3 dataset results, SMORE is best at $3\|0$, though in this case it is not a clear winner. The upper quartile appears to outperform ECLARE, but the mean is lower. These results are consistent across all three metrics. At all resolutions worse than $3\|0$, ECLARE has the best performance among all algorithms.
5 Discussion and conclusions
In this paper we presented ECLARE, a novel self-SR algorithm for anisotropic volumes that trains solely on data extracted from in-plane slices and thus requires no external training data. We presented a model of the imaging process in 2D multi-slice MRI and used ESPRESO [30] to estimate the slice selection profile. With an estimated slice profile, slice thickness and slice separation are both modeled, allowing for training data simulation from the HR in-plane data already present in the anisotropic volume. ECLARE uses a neural network architecture based on WDSR to efficiently train weights from scratch in a short amount of time, taking approximately 5 minutes on a single NVIDIA V100 GPU with a memory footprint under 2 GB. The result is a SR technique that respects the LR image contrast and presence of pathology, does not require external training data, and produces HR image volumes with the correct digital resolution and FOV. When compared to four other approaches on two simulated datasets, our approach achieved superior quantitative and qualitative results with respect to signal recovery metrics PSNR and SSIM as well as a downstream task metric—cDSC for brain region and MS lesion segmentation. This is true for all investigated resolutions except for $3\|0$, where SMORE performs best. This is likely due to two reasons: 1) SMORE does not have a model of slice gap, and therefore will perform best when no gap is present in the data; and 2) SMORE relies on pre-trained weights and was primarily tested and evaluated using data with resolutions $2\|0$ and $3\|0$.
We note that computation time is a primary limitation of ECLARE. ECLARE must fit a model for each multi-slice 2D MR volume. Although its training time is relatively fast compared to other self-SR methods, it is slower than pre-trained models that perform inference only. It would be possible to pre-train-then-fine-tune ECLARE as SMORE, but this introduces a tradeoff between time and relying on pre-training as a prior more than the in-plane data itself. Future work will investigate these tradeoffs in detail.
The SR of medical images has been a research area for decades. Data-driven approaches now dominate the field of SR, yet the problem is far from solved. When the test-time image does not belong to the training domain, performance of trained approaches is unpredictable. In medical imaging, domain issues may occur across a number of axes, including resolution differences, anatomical variations, presence of pathology, differences in contrast or scan acquisition parameters, noise level and distribution mismatch, or others. Our proposed method ECLARE considers all of these by training on in-plane data and leveraging a forward model inspired by the MR acquisition physics. Because it does not rely on external data for prior information, ECLARE has the potential to be used on any volumetric image and is not limited to scenarios for which a large amount of data—matched in both contrast and resolution—can be used for training.
Disclosures
This research is based on technology described in U.S. Patent No. $12,067,698$, titled ”Machine learning processing of contiguous slice image data,” assigned to Jerry L. Prince, Can Zhao, and Aaron Carass.
Acknowledgments
This material is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1746891 (Remedios). Development is partially supported by NIH ORIP grant R21 OD030163 (Pham). This work also received support from National Multiple Sclerosis Society RG-1907-34570 (Pham), FG-2008-36966 (Dewey), CDMRP W81XWH2010912 (Prince), NIH K01EB032898 (Schilling), and the Department of Defense in the Center for Neuroscience and Regenerative Medicine. The opinions and assertions expressed herein are those of the authors and do not reflect the official policy or position of the Uniformed Services University of the Health Sciences or the Department of Defense.